# A Class of Generalized Shift-Splitting Preconditioners for Double Saddle Point Problems

### v3

@@ 7 Numerical experiments
(f) $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B}$
(g) $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$
(h) $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$
@@
Figure 1: Eigenvalue distributions of $\mathcal{B},$ $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ for def__\__set.pow $=5$ with $\bm{\beta}=0.1$ for Example 1.
Parameter selection: For the DS preconditioner, we choose the parameter $\alpha$ (denoted by $\alpha_{\text{DS}}$) as follow [8]:
For the RDF preconditioner, the parameter $\alpha$ is selected from the interval $(0,1)$ with a step size of 0.01. The optimal performance, in terms of minimal CPU times, is achieved with smaller values of $\alpha$ as found in [11]. For the SS preconditioner, we take $\alpha=0.01.$ For the GSS, RGSS-I and RGSS-II preconditioners parameters are selected as follows: $\alpha_{\text{DS}}=\beta=0.01,$ $\tau=0.001,$ $P=A,$ $Q=CC^{T}$ and $R=I.$ The optimal parameter $\bm{\omega}$ (denoted by $\bm{\omega}_{exp}$) is determined experimentally within the interval $[2,30]$ with step size one, which yields minimal CPU times.
Numerical results: The numerical results for GMRES and various PGMRES methods with $\bm{\beta}=0.1$ and $0.001$ are presented in Tables 1 and 2. We observe that the GMRES method exhibits a significantly slower convergence rate compared to all other PGMRES methods, even does not converge within $5000$ iterations when def_setup.pow $=7$. On the other hand, we observe that our proposed preconditioners outperform all the compared preconditioners in terms of both IT and CPU times. For the DS and SS preconditioners, the IT increases as the size of the saddle point matrix $\mathcal{B}$ grows. Whereas the proposed GSS, RGSS-I and RGSS-II preconditioners maintain a consistent IT regardless of matrix size.
Eigenvalue distributions: In order to better illustrate the superiority of the proposed GSS, RGSS-I and RGSS-II preconditioners, the eigenvalue distribution of $\mathcal{B}$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ are displayed in Figure 1. From Figure 1, we observe that the eigenvalues of the preconditioned matrices $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B},$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ have clustered better than the coefficient matrix $\mathcal{B},$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B}.$ This indicates enhanced computational efficiency, highlighting the superiority of the GSS, RGSS-I, and RGSS-II preconditioners over existing methods.
Influence of the parameters ${\alpha,\beta,\tau,\bm{\omega}}$:
(a) By varying $\alpha=\beta$ within the interval $[0.01,1]$ and $\omega$ within the range $[1,30]$ for the GSS preconditioner.
Figure 2: Convergence curves of the GSS and RGSS-II RGMRES methods varying the parameters $\alpha,$ $\beta,$ $\tau,$ $\bm{\omega}$ for Example 1 with $\bm{\beta}=0.1$.
Condition number analysis:
Figure 3: Relationship between condition numbers of the preconditioned matrices $\mathcal{M}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ varying the parameter $\bm{\omega}$ in $[1,100]$ with $\bm{\beta}=0.1$ for Example 1.
To evaluate the robustness of the proposed GSS, RGSS-I and RGSS-II preconditioners, we assess the condition numbers of the preconditioned matrices $\mathcal{M}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$. For any nonsingular matrix $A$, the condition number $\kappa(A)$ is defined as $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2},$ where $\|A\|_{2}$ denotes the spectral norm of $A.$ The condition number of $\mathcal{B}$ is $053.6396e+$ 05 which is very large, making the system (1.1) ill-conditioned to solve. In Figure 3, we depict the effect of the parameter $\bm{\omega}$ ranges from 1 to 100 with a step size of one on the preconditioned matrices $\mathcal{M}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ for the case def_setup.pow $=5$. We observe that for all values of $\bm{\omega}$, the condition numbers of the preconditioned matrices remain within the range $[1,1.5]$. This indicates that the preconditioned systems are well-conditioned, demonstrating that the GSS, RGSS-I, and RGSS-II preconditioners are robust and effective.
Analyzing the sensitivity of the solution:
Figure 4: Behaviour of the relative errors in the solution with with increasing $n_{p}$ for def_setup.pow $=5,6,7$ with $\bm{\beta}=0.1$ for Example 1.

### v4

@@ 7 Numerical experiments
(a) $\mathcal{B}$
(b) $\mathcal{P}_{\text{BD}}^{-1}\mathcal{B}$
(c) $\mathcal{P}_{\text{BP}}^{-1}\mathcal{B}$
(d) $\mathcal{P}_{\text{DS}}^{-1}\mathcal{B}$
(e) $\mathcal{P}_{\text{RDF}}^{-1}\mathcal{B}$
(f) $\mathcal{P}_{\text{SS}}^{-1}\mathcal{B}$
(g) $\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B}$ $(\bm{\omega}=30)$
(h) $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ $(\bm{\omega}=30)$
(i) $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ $(\bm{\omega}=26)$
Figure 1: Eigenvalue distributions of $\mathcal{B},$ $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{BP}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ for def__\__set.pow $=5$ with ${\bm{\nu}}=0.1$ for Example 1.
The BD preconditioner is utilized in its exact form, as specified in [13, equation $(5.6)$]. We adopt the BP preconditioner as defined in [26, equation $(2.1)$]. The DS preconditioner requires solving two linear subsystems with coefficient matrices $(\alpha I+A+\frac{1}{\alpha}B^{T}B)$ and $(\alpha I+A+\frac{1}{\alpha}CC^{T})$. Similarly, the RDF preconditioner involves two linear subsystems with coefficient matrices $(A+\frac{1}{\alpha}B^{T}B)$ and $(A+\frac{1}{\alpha}CC^{T})$. The SS preconditioner necessitates solving linear subsystems with coefficient matrices $(\alpha I+2\bm{\nu}M)$, $(\alpha I+M)$, and $(\alpha I+M(\alpha I+2\bm{\nu}M)^{-1})M+K(\alpha I+M)K^{T}$. For all these subsystems, we employ exact solvers, such as Cholesky factorization. Furthermore, in Algorithms 1, 2, and 3, all linear subsystems are solved using Cholesky factorization. For the inexact versions of our proposed preconditioners $\widetilde{\mathcal{P}}_{\mathrm{GSS}},$ $\widetilde{\mathcal{P}}_{\text{RGSS-I}}$ and $\widetilde{\mathcal{P}}_{\text{RGSS-II}}$ (abbreviated as “IGSS”, “IRGSS-I” and “IRGSS-II”, respectively), we construct $\widetilde{P},\,\widetilde{P}_{1},\,\widetilde{Q},$ and $\widetilde{Q}_{1}$ as follows: consider $\widetilde{A}=L_{A}L^{T}_{A},\,\widetilde{A}_{p}=L_{A_{p}}L^{T}_{A_{p}},\widetilde{D}=L_{D}L^{T}_{D},\,\widetilde{D}_{q}=L_{D_{q}}L^{T}_{D_{q}},$ $\widetilde{P}=\mathrm{diag}(B\widetilde{A}_{p}^{-1}B^{T}),$ $\widetilde{P}_{1}=\mathrm{diag}(B\widetilde{A}^{-1}B^{T}),$ $\widetilde{Q}=\mathrm{diag}(C^{T}\widetilde{D}_{q}^{-1}C),$ and $\widetilde{Q}_{1}=\mathrm{diag}(C^{T}\widetilde{D}^{-1}C),$ where $L_{A},\,L_{A_{p}}\,L_{D},$ and $L_{D_{q}}$ are the incomplete Cholesky factorization of $A,\,(\alpha P+\bm{\omega}A),\,D,$ and $(\beta Q+\bm{\omega}D),$ respectively, produced by the Matlab function:
[…]
Parameter selection: For the DS preconditioner, we choose the parameter $\alpha$ (denoted by $\alpha_{\text{DS}}$) as follows [8]:
For the RDF preconditioner, the parameter $\alpha$ is selected from the interval $(0,1)$ with a step size of 0.01. The optimal performance, in terms of minimal CPU times, is achieved with smaller values of $\alpha$ as found in [11]. For the SS preconditioner, we take $\alpha=0.01.$ For the GSS, RGSS-I, and RGSS-II preconditioners, as well as for their inexact forms, the parameters are chosen as follows: $\alpha=\beta=0.01,$ $\tau=0.001,$ $P=A,$ $Q=CC^{T},$ and $R=I.$ The optimal parameter $\bm{\omega}$ (denoted by $\bm{\omega}_{exp}$) is determined experimentally within the interval $[2,30]$ with step size one, which yields minimal CPU times.
Numerical results: The numerical results for GMRES and various PGMRES methods with ${\bm{\nu}}=0.1$ and $0.001$ are presented in Tables 1 and 2. We observe that the GMRES method exhibits a significantly slower convergence rate compared to all other PGMRES methods, even does not converge within $5000$ iterations when def_setup.pow $=7$. On the other hand, we observe that our proposed preconditioners outperform all the compared preconditioners in terms of both IT and CPU times. For the DS and SS preconditioners, the IT increases as the size of the saddle point matrix $\mathcal{B}$ grows. Whereas the proposed GSS, RGSS-I and RGSS-II preconditioners maintain a consistent IT regardless of matrix size. Moreover, from Tables 1 and 2, we observe that the inexact preconditioners IGSS, IRGSS-I, and IRGSS-II require less CPU time to converge, demonstrating more efficiency compared to their exact counterparts GSS, RGSS-I, and RGSS-II, respectively.
Eigenvalue distributions: In order to better illustrate the effectiveness of the proposed GSS, RGSS-I and RGSS-II preconditioners, the eigenvalue distribution of $\mathcal{B}$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{BP}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B}\,{(\bm{\omega}=30)},$ $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}\,{(\bm{\omega}=30)}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}\,{(\bm{\omega}=26)}$ (with ${\bm{\nu}}=0.1$) are displayed in Figure 1. From Figure 1, we observe that the eigenvalues of the preconditioned matrices $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B},$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ have clustered better than the coefficient matrix $\mathcal{B},$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{BP}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B}.$ The spectral radius of the saddle point matrix $\mathcal{B}$ is $\vartheta(\mathcal{B})=3.9872$. The eigenvalues of $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B}$ are all real with $\vartheta(\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B})=1.8019$ and $\vartheta({\mathcal{P}}_{\text{BP}}^{-1}\mathcal{B})=1.0128.$ The preconditioned matrices $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B}$ both have a spectral radius of 1. However, $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B}$ has non-real eigenvalues with imaginary parts of order $10^{-1}$ or $10^{-2}$, whereas the imaginary parts of the non-real eigenvalues of $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B}$ are significantly smaller, on the order of $10^{-8}$. The computed spectral radius for preconditioned matrices $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ is $\vartheta(\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B})=1.9995$, $\vartheta(\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B})=0.0333$, $\vartheta(\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B})=0.0333$, and $\vartheta(\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B})=0.0385,$ respectively. However, the complex part of the non-real eigenvalues of the SS preconditioned matrix is in the range of $10^{-1},$ whereas for GSS and RGSS-I, they are in the range of $10^{-5}.$ There are at least $961(=n)$ eigenvalues of RGSS-I preconditioned matrix is equal to $0.0333$ and for the RGSS-II preconditioned matrix, at least $1922(=n+l)$ real eigenvalues are equal to $0.0385$ Thus, we can conclude that GSS, RGSS-I, and RGSS-II preconditioned matrices have better clustered spectrums compared to other preconditioned matrices. This indicates enhanced computational efficiency, highlighting the effectiveness of the GSS, RGSS-I, and RGSS-II preconditioners.
Influence of the parameters ${\alpha,\beta,\tau,\bm{\omega}}$:
(a) By varying $\alpha=\beta$ within the interval $[0.01,1]$ and $\omega$ within the range $[1,30]$ for the GSS preconditioner
Figure 2: Convergence curves of the GSS and RGSS-II RGMRES methods varying the parameters $\alpha,$ $\beta,$ $\tau,$ $\bm{\omega}$ for Example 1 with ${\bm{\nu}}=0.1$.
Condition number analysis:
Figure 3: Relationship between condition numbers of the preconditioned matrices $\mathcal{P}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ varying the parameter $\bm{\omega}$ in $[1,100]$ with ${\bm{\nu}}=0.1$ for Example 1.
To evaluate the robustness of the proposed GSS, RGSS-I, and RGSS-II preconditioners, we assess the condition numbers of the preconditioned matrices $\mathcal{P}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$. For any nonsingular matrix $A$, the condition number $\kappa(A)$ is defined as $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2},$ where $\|A\|_{2}$ denotes the spectral norm of $A.$ The condition number of $\mathcal{B}$ is $053.6396e+$ 05, which is moderately hight. On the other hand, in Figure 3, we depict the effect of the parameter $\bm{\omega}$ ranges from 1 to 100 with a step size of one on the preconditioned matrices $\mathcal{P}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ for the case def_setup.pow $=5$. We observe that for all values of $\bm{\omega}$, the condition numbers of the preconditioned matrices remain within the range $[1,1.5]$. This indicates that the preconditioned systems are well-conditioned, demonstrating that the GSS, RGSS-I, and RGSS-II preconditioners are robust and effective.
Analyzing the sensitivity of the solution:
Figure 4: Behaviour of the relative errors in the solution with with increasing $n_{p}$ for def_setup.pow $=5,6,7$ with ${\bm{\nu}}=0.1$ for Example 1.
Example 2.
The leaky lid driven cavity problem [17, 18]: We consider the DSPP arising from the discretization of the incompressible Stokes flow problem, which is given by
$\begin{array}[]{rl}-\Delta{\bm{u}}+\nabla\bm{v}=\bm{0},&\leavevmode\nobreak\ \text{in}\quad\Xi,\\ \quad\quad\quad\nabla\cdot{\bm{u}}=0,&\leavevmode\nobreak\ \text{in}\quad\Xi,\\ \quad\quad\quad\bm{u}=\bm{0},&\leavevmode\nobreak\ \text{on}\quad\Xi\backslash\partial\Xi_{lid},\\ \quad\quad\quad\bm{u}_{x}=1,&\leavevmode\nobreak\ \text{on}\quad\partial\Xi_{lid},\end{array}$ (7.10)
[…]
Table 3: Experimental results of GMRES, BD, BD-MINRES, DS, RDF, SS, GSS, RGSS-I and RGSS-II PGMRES methods for Example 2.
Process Grids $16\times 16$ $32\times 32$ $64\times 64$ $128\times 128$
$\mathcal{B}$) $770\times 770$ $2946\times 2946$ $11522\times 11522$ $45570\times 45570$
GMRES IT $246$ $422$ $692$ $1169$
CPU $0.2042$ $2.7629$ $158.8824$ $1880.7285$
BD IT $38$ $38$ $38$ $38$
CPU $0.5025$ $5.2317$ $104.2237$ $1067.3401$
BD-MINRES IT $32$ $32$ $32$ $32$
CPU $0.3213$ $4.5613$ $90.6470$ $650.6753$
DS IT $30$ $39$ $49$ $59$
CPU $0.5662$ $3.6632$ $104.0399$ $885.1938$
RDF IT $11$ $13$ $12$ $11$
CPU $0.4619$ $1.5057$ $26.8306$ $501.8811$
SS IT $11$ $18$ $32$ $53$
CPU $0.3426$ $1.9709$ $48.6013$ $1675.7429$
GSS IT $2$ $2$ $2$ $3$
$\bm{\omega}_{exp}=25$ CPU $0.2040$ $0.7120$ $6.0955$ $205.7726$
RGSS-I IT $2$ $2$ $2$ $3$
$\bm{\omega}_{exp}=29$ CPU $0.2296$ $0.6767$ $5.3381$ $207.4075$
RGSS-II IT $2$ $2$ $2$ $3$
$\bm{\omega}_{exp}=29$ CPU $0.2199$ $0.6125$ $5.4854$ $209.7603$
Discretizing the systems in (7.10) using the IFISS software package [17] with $Q2-Q1$ finite elements on uniform leads to the DSPP (1.1). The right hand side vector $\bm{b}\in{\mathbb{R}}^{n+l+m}$ is chosen so that the exact solution of the DSPP is $\bm{u}=[1,1,\ldots,1]\in{\mathbb{R}}^{n+l+m}.$ We compare our proposed GSS, RGSS-I and RGSS-II PGMRES methods with the GMRES method and PGMRES methods with BD, BD-MINRES, DS, RDF, and SS preconditioners, respectively. The BD preconditioner is used in its exact form as given in [13, equation (6)]. The implementation of all the preconditioners is done as in Example 1. We consider the grids of sizes $16\times 16$, $32\times 32$, $64\times 64$, and $128\times 128$
Parameter selection: The parameter choices for the tested preconditioner are taken as in Example 1. In addition, for the proposed preconditioners, we consider $\tau=0.0001.$ The optimal parameter $\bm{\omega}$ (denoted as $\bm{\omega}_{\text{exp}}$) is experimentally determined by testing values in the interval $[10,30]$ with a step size of one, selecting the value that minimizes CPU time.
@@
(f) $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B}$ $(\bm{\omega}=25)$
(g) $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ $(\bm{\omega}=29)$
(h) $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ $(\bm{\omega}=29)$
@@
Figure 5: Eigenvalue distributions of $\mathcal{B},$ $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ for Example 2 with grid $16\times 16$.
Numerical results and discussions: Table 3 presents the numerical results for GMRES and various PGMRES methods. The standard GMRES method converges significantly slower than all PGMRES variants, requiring a substantially larger number of IT. In contrast, our proposed preconditioners outperform the compared ones in IT and CPU time. Notably, they achieve convergence within just 2 or 3 iterations, highlighting the efficiency of the GSS, RGSS-I, and RGSS-II preconditioners compared to the baseline preconditioners.
Eigenvalue distributions: To better illustrate the effectiveness of the proposed preconditioners, Figure 5 presents the eigenvalue distributions of $\mathcal{B}$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\text{GSS}}^{-1}\mathcal{B}\,{(\bm{\omega}=25)},$ $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}\,{(\bm{\omega}=29)}$ and $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}\,{(\bm{\omega}=29)}$ with ${\bm{\nu}}=0.1$ are displayed in Figure 5. Moreover, we computed the spectral radius of $\mathcal{B}$ and each preconditioned matrices, which are $\vartheta(\mathcal{B})=7.6198,$ $\vartheta(\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B})=1.8018,$ $\vartheta(\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B})=0.9997,$ $\vartheta(\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B})=1,$ $\vartheta(\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B})=1.9974$, $\vartheta(\mathcal{P}_{\mathrm{GSS}}^{-1}\mathcal{B})=0.04$, $\vartheta(\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B})=0.0345$, and $\vartheta(\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B})=0.0345.$ Therefore, based on the spectral radius values and observation from Figure 5, it leads to the conclusion that the GSS, RGSS-I, and RGSS-II preconditioned matrices exhibit a more tightly clustered spectrum compared to other preconditioned matrices. This highlights the enhanced effectiveness of our proposed preconditioners.